OpenStudy (anonymous):

Solution of x^4+1=0 in real numbers

OpenStudy (shubhamsrg):

convert into quadratic and see for yourself if solution is possible .. i.e. let x^2 = y then y^2 +1 =0 easier to tell now?? well was easier earlier also..

mathslover (mathslover):

transpose 1 to RHS first

mathslover (mathslover):

what do you get?

mathslover (mathslover):

@wazim please reply

OpenStudy (anonymous):

x^4+1=0 x^4=-1 Since there is no real number that will give -1 its power is raised to 4.

OpenStudy (anonymous):

So what do u think?

OpenStudy (anonymous):

i didnt get what is meant to take transpose of equatiom

mathslover (mathslover):

transpose means like this: \[\large{x^4+1=0}\] \[\large{x^4+1-1=0-1}\]

mathslover (mathslover):

can you tell me what will u get as a simplified form?

mathslover (mathslover):

transpose 1 to RHS means : subtract 1 both sides what do you get @wazim ?

OpenStudy (anonymous):

\[x^{4}-1=-2\]

mathslover (mathslover):

what is 1-1 ?

OpenStudy (anonymous):

oh yes its zero. and x^4 =-1. so no solution in reals

mathslover (mathslover):

right : \[\large{x^4=-1}\] \[\large{\sqrt[4]{x^4}=\sqrt[4]{-1}}\]

mathslover (mathslover):

what do you get now @wazim ?

OpenStudy (anonymous):

Complex number

mathslover (mathslover):

x = ?

mathslover (mathslover):

\[\large{x=(-1)^{\frac{1}{4}}}\] \[\large{x=((-1)^{\frac{1}{2}})^{\frac{1}{2})}}\] \[\large{x=i^{\frac{1}{2}}}\] \[\large{x=\sqrt{i}}\] M i right?

OpenStudy (anonymous):

YES

mathslover (mathslover):

so that is ur answer

OpenStudy (anonymous):

can we resolve it into partial fraction \[\frac{ 1 }{ x^{4}+1 }\]